# Effective Potential and Topological Photon Spheres: A Novel Approach to Black Hole Parameter Classification

### v1

@@ Abstract
In this article, we anchor our analysis on the premise that the presence of a photon sphere is an intrinsic characteristic of any ultra-compact gravitational structure with spherical symmetry. Utilizing the concept of a topological photon sphere, we categorize the behaviors of diverse gravitational models based on the structure of their photon spheres. This innovative approach enables us to define boundaries for black hole parameters, which subsequently allows us to classify the model as either a black hole or a naked singularity. Indeed, we will demonstrate that the existence of this reciprocal relationship between the gravitational structure and the presence of a photon sphere offers a unique advantage that can be directly and inversely harnessed. Furthermore, we present a precise correlation between the behavior of the defined effective potential and the total topological charge for each model. Through an independent examination of the effective potential, we show that a gravitational model typically exhibits naked singularity behavior when at least one effective potential minimum (a stable photon sphere) protrudes beyond the event horizon. Finally, we will compare the effect of adding perfect fluid dark matter to the models under study and observe the impact of this quantity on the behavior of the photon sphere and the extent of their dominance.
@@ 1 Introduction
In the study of ultra-compact gravitational objects, the simplest and initial step for distinguishing between black hole configurations and naked singularities is the examination of the metric function followed by the identification of the event horizon. The emergence of a metric function is, in fact, the result of a complex amalgamation of initial conditions such as energy conditions and geometric compatibility with the equations of general relativity, among others, which ultimately manifest themselves in the form of a mathematical-physical metric function for the models. The remaining weaknesses of the resulting models, are then covered by conjectures such as the Weak Cosmic Censorship Conjecture (WCCC) through the creation of an event horizon. This conjecture posits that singularities that appear while solving Einstein’s equations must be concealed behind an event horizon. This is because if singularities were observable from the rest of space-time, causality would be jeopardized, and physics would lose its predictive power. It seems that this condition, rather than being a powerful scientific statement, is an attempt to cover the current scientific and instrumental weaknesses and our lack of understanding of the geometric spacetime that we are, in fact, unable to describe.
It should be noted that the non-fulfillment of the WCCC condition does not imply ’non-existence’ or ’absence’ but rather the ’existence’ of a space-time which, although it has its own gravitational effects, our current information is not enough to decipher this space, and even today, extensive studies are being conducted on the possibility and conditions of the emergence and observation of these regions. For example, space-time singularities created during gravitational collapse can be observed by an external observer, and the visibility of a spacetime singularity depends on the initial conditions of the collapsing matter [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Therefore, knowing under what conditions our model behaves like a black hole and from what range it transitions to an ultra-compact object without an event horizon (naked singularity) is important. But the regions of the naked singularity may also be involved in the results under certain conditions. Hence, the extent of their influence must also be defined.
After constructing the initial metric function based on effective theories in action, what will most influence the location of the event horizon are the values considered for the parameters related to each theory. In simple terms, these parameters’ permissible range can be studied based on the roots of the metric function. However, it is clear that this method will definitely not be applicable in determining the dominance range of the naked singularity.
@@
So, in order to gain a better understanding of the overall space surrounding an ultra-compact object, one must seek a potential function that provides a more comprehensive answer to how the gravity of this structure affects the entire surrounding space. In pursuit of such a function, we know on one hand that the nature of attractive structures, whether in their classical or quantum form, necessitates the existence of stable and unstable circulating loops. On the other hand, in gravity, the presence of such loops around a black hole or an ultra-compact object has been well calculated both experimentally and observationally, as well as theoretically for different black hole models. In the observational study, investigators documented the presence of a distinct dark shadow, which emerged due to the gravitational deflection of light by the colossal black hole located at the nucleus of the massive elliptical galaxy M87. This phenomenon conclusively substantiates the existence of photon rings encircling the central dense object. [15, 16, 17, 18, 19, 20] or, theoretically, Cunha et al. showed during their studies that, in spherical symmetry, standard black holes, as well as other ultra-compact objects (that can be with or without event horizon) in general relativity, can have planar circular photon orbits. In such a way that the stable type causes instability, and its unstable type could determine the shadows of the black hole [21]. Also, in a new paper it was shown in a geometrical proof that ’any spherically symmetric space-time that is asymptotically flat and has a horizon must contain a photon sphere, a spherical photon shell’ [22]. Therefore, considering the above characteristics, it seems that the idea of the photon sphere can be used for a more precise classification of the space surrounding the ultra-compact structures.
To calculate the photon sphere in the usual way, one must go to the action and, after calculating the Lagrangian and Hamiltonian, extract the effective potential and then study the photon ring. But, in this work, instead of using the traditional method above for calculating the photon ring, which especially with the complication of the action requires more complex mathematical calculations, we propose the use of topological photon sphere. Of course, it must be stated that the use of topological photon spheres, in addition to fulfilling our need for spatial classification, will also bring other valuable benefits for us, which will be referred to at the appropriate time.
@@
As, we have stated, the foundation of our work in this article is predicated on the necessity of the photon sphere for ultra-compact object structures. However, to approach the topological photon sphere, we commence our study of the photon sphere from the point where Cardoso and his colleagues have articulated: ”The mere observation of an unstable light ring is strong evidence for the existence of black holes”, or, ”The existence of a stable light ring is thus an unavoidable feature of any ultra-compact star” [23]. Then Cunha et al. showed during their studies that, in spherical symmetry, standard black holes, as well as other ultra-compact objects (that can be with or without event horizon) in general relativity, can have planar circular photon orbits. In such a way that the stable type causes instability, and its unstable type could determine the shadows of the black hole [22]. Finally, after Cunha et al. which extended their work to spherically symmetric 4D black holes[26],” Shao-Wen Wei” used this idea and” Duan” mapping, and extended the discussion of the photon ring to photon sphere with a topological approach. He stated that there exists at least one standard photon sphere outside the black hole not only in asymptotically flat space-time, but also in asymptotically AdS and dS space-time [24].
@@ 3 Explanation of topological charges based on properties of effective potential
In his paper, ’Wei’ demonstrated through the examination of winding numbers and the calculation of total charges for a model that whenever the TTC is -1, the structure will exhibit black hole behavior, and conversely, if the TTC is 0, the structure will be in the form of a naked singularity [24]. We have extended this work to show that in addition to the value of 0, the TTC value of +1 can also appear as a naked singularity[25]. However, in this section, we do not wish to delve into the meaning of the values mentioned above with regard to the field configuration around zero points and the analysis of winding numbers accordingly. Instead, we want to examine and study the concept of each of the values 0,+1, and -1 from the perspective of the effective potential equation (15). The potential defined in equation (15) for the study of the topological photon sphere, being a potential structure in its entirety, allows us to expect that classical principles governing potential functions can be utilized for its interpretation. This means that the extremum points on the potential curve play a significant role in justifying the physical behavior of the system dominated by the potential. If the potential curve has a local or global maximum, the system will leave this maximum upon the slightest perturbation. Conversely, if the system has a local minimum, it will tend to maintain its state unless there is a stronger global minimum in its vicinity, in which case, upon reaching the necessary disturbance, the system will move towards the global minimum for greater stability. Here, we will demonstrate that this same behavioral pattern holds true for our topological potential as well.
@@
To understand why, from an energy perspective, TTC = -1 corresponds to an unstable photon sphere, we have depicted a scenario for the Hayward black hole in Fig (1). As observed in Fig (1a), this black hole possesses a single zero point outside its event horizon, resulting in a TTC of -1, in accordance with reference [24]. In Fig (1b), we have plotted the potential H(r). It is evident that this potential has a global maximum, which upon closer inspection of Fig (1c), is precisely located at the zero point or in other words, at the same location as the photon sphere. Given that this point represents a global maximum, any photon trapped within this potential will abandon this peak, or more precisely, the generating factor of this maximal potential (our black hole), at the slightest disturbance in pursuit of a state with lower energy. This signifies that we possess an unstable photon sphere, precisely the phenomenon of observational interest.
@@
From the energy perspective, these two states occur when one or several local or global minima appear outside the event horizon in such a way that they fully influence the nearby maxima. For a better understanding of the above concept, we examine these conditions for the Hayward model at a TTC of 0 and the Hyperscaling violation model [25] at a TTC of +1.
In a simplified assumption, devoid of complexities, if we consider the attraction in the form of potential function minima, as evident in Fig (2), it appears that in the Hayward model, an increase in the magnetic charge (g) aligns with gravity to enhance the gravitational strength of the model by creating local minima outside the event horizon. As we observe in Fig (2b), the presence of a minimum - a stable photon sphere - at r = 1.389900878 in the vicinity of a maximum - an unstable photon sphere - implies that even if a photon could leave the maximum under some perturbation, it would most likely be trapped in the nearest minimum, which would be our stable photon sphere. Two points must be noted here. First, this typically occurs when, from the metric function perspective, we are in a rootless region or what is known as a naked singularity space. In other words, the emergence of potential minima beyond the event horizon, which according to our assumption somehow confirms the extension of the model’s intense gravity domain, becomes synonymous with a naked singularity. The second point to consider is how far increasing the parameters of the black hole (here g) will expand the dominance of stable photon spheres? In other words, does the naked singularity have its limits? In response, it should be said that, for example, in the Hayward model, the increase in the magnetic parameter is limited by the initial parameter values such as mass and de Sitter radius, and for our choices (m = 1, l = 1), the parameter (g) will only be responsive up to the value of (1.21834146), and beyond this range, neither the metric function nor the potential function will practically have a response for space, which we refer to as the unauthorized region.
Figure 1: Fig (1a):The normal vector field $n$ in the $(r-\theta)$ plane. The photon sphere is located at $(r,\theta)=(3.002309062,1.57)$ with respect to $(g=-0.21834,m=1,l=1)$ , (1b): the topological potential H(r) for regular Hayward AdS black hole model , (1c): Enlarged details of diagram (1b)
@@
In Fig(3), we encounter a scenario with a total topological charge (TTC) of +1. The fundamental difference between this state and the state with TTC = 0 is that here, due to the intensity of gravity, either no maximum is formed at all, or if an unstable photon sphere does form, as in Fig (3c) at (r = 1.879279979), the sheer number of local minima in the vicinity is so great that the possibility of escape, even for photons, becomes unfeasible. Considering that if photons cannot escape such an environment, it can be inferred that no other massive particles would be able to escape either, effectively evoking conditions akin to those within an event horizon. Therefore, the emergence of a minimum in the effective potential outside the event horizon can be interpreted as a loss of all spatial information, and consequently, a violation of the Weak Cosmic Censorship Conjecture (WCCC), which equates to the manifestation of a naked singularity. However, we emphasize that the loss of information and spatial geometry does not imply non-existence, as the structure continues to exhibit gravitational evidence, namely the photon sphere, indicating its ongoing presence and vitality.
@@ 4 Effective potential and topological photon sphere as a tool for classification
Initially, it is prudent to briefly address the methodology of this study, along with its advantages and disadvantages, before proceeding to the selected models. In this work, we begin by constructing a potential function (15) utilizing the metric function of the model, which necessarily must be spherically symmetric. Based on this, we will construct the vector field $\phi$ (16). Subsequently, on the equatorial plane ($\theta=\frac{\pi}{2}$), we locate the roots of this vector field, which, as previously mentioned, will be the sites of the photon spheres. We then study the topological charge of these locations.
For the first time, based on the TTC and aligning its results with those of the effective potential graph, as well as the initial parameter values of the model and the roots of the metric function, we classify the space surrounding an ultra-compact object. We will also determine, with relatively good precision, the range of variation for the specific parameter of the model.
It must be acknowledged that the utilization of this method carries its own set of advantages and disadvantages. The traditional approach to studying photon spheres necessitates the derivation of the Lagrangian from the action, followed by the construction of the Hamiltonian. Once the Hamiltonian is obtained, one must determine the effective potential and then proceed to study the photon sphere, taking into account the symmetries of the Killing vectors and their commutativity with respect to the model’s symmetries. It is evident that the addition of various theories to the action can significantly complicate this pathway, both mathematically and computationally.
@@
One of the significant advantages of this method is the direct use of the metric function to construct the potential, bypassing the aforementioned concerns. Utilizing the equatorial plane for the vector field ($\phi$), analogous to the Poincaré plane, and the resultant dimensional reduction is another notable benefit. Ultimately, the ability to study a broader scope of the space surrounding an ultra-compact object is a fundamental point that leads to the classification of this space based on this method. Finally, a practical concept can also be introduced using this method, which we called the possible radius limit for the photon sphere.
@@
$R_{PLPS}$:According to the allowed range of parameters to preserve the black hole structure, we called the minimum or maximum possible radius for the appearance of an unstable photon sphere as the $R_{PLPS}$.
@@
However, there are essential drawbacks to consider. Since the vector field ($\phi$) needs to be separable with respect to spatial coordinates, only diagonal metrics with spherical symmetry can be employed. Given that most studied models possess a spherical structure, and the remainder can often be expressed in diagonal form using the C-metric, this limitation can be somewhat overlooked. Another issue is that since the Poincaré plane is used for dimensional reduction, this method may not be applicable to models with fewer than four dimensions. Finally, as this is a nascent approach, there may be weaknesses that will become apparent over time. As a point of interest at the conclusion of this section, it should be noted that the simplifications, computational power, and the novelty of this method and mathematical model for studying the symmetric potential functions have garnered such attention that the mathematical structure of the model has been extended to thermodynamics. Consequently, extensive studies based on the mathematical model of this method have been conducted on the phase transition of black hole models [30, 31, 32, 33, 34, 35, 36, 37, 38].
@@ 4.1 Photon Sphere and perfect fluid dark matter black hole(PFDM)
In the first case, considering the chosen value of the parameter $\lambda=0.62$ and the presence of the event horizon ($r_{h}=2.97163$) Fig (4), we observe the appearance of a photon sphere outside the event horizon. This photon sphere possesses a topological charge of -1 Fig (5a) and, as evident in Fig (5b), represents an energy maximum. Consequently, in this case, we are confronted with a black hole that contains an unstable photon sphere. However, in the second case, our choice of $\lambda=2.15$ results in a metric function without roots, which is clearly visible in Fig (4). The gravitational structure displays two photon sphere with charges of -1 and +1, indicating a space-time with a TTC of zero, Fig (6a). Energy-wise, this situation is equivalent to the emergence of both a minimum and a maximum, as shown in Fig (6b). Since the created minimum prevents even photons from escaping such a space, observationally, the space is completely dark. Moreover, since there is no event horizon, it can be assumed that we are dealing with conditions similar to those inside an event horizon. Given these conditions, in this scenario, the structure manifests itself in the form of a naked singularity.
Now we have to answer the question that for what values of the parameter $\lambda$ the space will be in the form of a black hole, and for any value outside this range, will our gravitational structure correspond to a naked singularity? We have shown the answers to these questions in Table 1.
@@
Here it is necessary to note that we have focused on the domain of $\lambda$ and assigned an arbitrary value to M. Obviously, changing these values can change our parameter range.

### v2

@@ Abstract
In this paper, we base our analysis on the assumption that the existence of a photon sphere is an intrinsic feature of any ultra-compact gravitational structure with spherical symmetry. Utilizing the concept of a topological photon sphere, we categorize the behaviors of various gravitational models based on the structure of their photon spheres. This innovative approach enables us to define boundaries for black hole parameters, subsequently allowing us to classify the model as either a black hole or a naked singularity. Indeed, we will demonstrate that the presence of this interplay between the gravitational structure and the existence of a photon sphere is a unique advantage that can be utilized from both perspectives. Our observations indicate that a gravitational model typically exhibits the behavior of a horizonless structure (or a naked singularity) when a minimum effective potential (a stable photon sphere) appears within the studied spacetime region. Additionally, in this study, we tried to investigate the effect of this structure on the behavior of the photon sphere by choosing models that are affected by the Perfect Fluid Dark Matter (PFDM). Finally, by analyzing a model with multiple event horizons, we show that the proposed method remains applicable even in such scenarios.
@@ 1 Introduction
In the study of ultra-compact gravitational objects, the simplest and most fundamental step to distinguish between different configurations is to examine the metric function and subsequently identify the event horizon. The metric function is essentially the result of integrating equations based on diverse and complex initial conditions (such as energy conditions, geometric compatibility with general relativity equations, etc.). Any remaining weaknesses that might challenge the physics of the derived model are addressed by conjectures such as the Weak Cosmic Censorship Conjecture (WCCC), which posits the creation of an event horizon. This conjecture assumes that singularities emerging from solving Einstein’s equations must be hidden behind an event horizon. This is because if singularities were observable from the rest of spacetime, causality would be compromised, and physics would lose its predictive power. It appears that this condition is more of an attempt to cover current scientific and instrumental weaknesses and our lack of understanding of geometric spacetime, which we are actually unable to describe. It is important to note that the failure to meet the WCCC condition does not imply ”non-existence” or ”absence,” but rather the ”existence” of a spacetime that, although it has its own gravitational effects, our current knowledge is insufficient to decode it. Extensive studies are currently being conducted on the possibility and conditions for the emergence and observation of these regions. For instance, spacetime singularities formed during gravitational collapse can be observed by an external observer, and the observability of a spacetime singularity depends on the initial conditions of the collapsing matter [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Therefore, it is crucial to know under what conditions our model behaves like a black hole and from what range it transitions to an ultra-compact object without an event horizon (naked singularity). However, naked singularity regions may also be involved under certain conditions in the results. Hence, their extent of influence must also be determined.
@@
After constructing the initial metric function, the most significant factor influencing the location of the event horizon is the values assigned to the parameters of each theory used in constructing the metric function. In a simplified view, the permissible range of these parameters can be studied based on the roots of the metric function. However, it is clear that this method will not be applicable in determining the dominance range of a naked singularity. Therefore, to gain a better understanding of the overall space surrounding an ultra-compact object, one must seek a potential function that provides a more comprehensive response to how the gravity of this structure affects the entire surrounding space. In pursuit of such a function, we know that the nature of gravitational structures, whether in their classical or quantum forms, necessitates the existence of stable and unstable circular orbits. Moreover, in gravity, the existence of such orbits around a black hole or an ultra-compact object has been well calculated both experimentally and observationally, as well as theoretically for various black hole models.For example, in a observational study, investigators documented the presence of a distinct dark shadow, which emerged due to the gravitational deflection of light by the colossal black hole located at the nucleus of the massive elliptical galaxy M87. This phenomenon conclusively substantiates the existence of photon rings encircling the central dense object [15, 16, 17, 18, 19, 20]. Also, theoretically, Cunha et al. showed during their studies that, in spherical symmetry, standard black holes, as well as other ultra-compact objects (that can be with or without event horizon) in general relativity, can have planar circular photon orbits. In such a way that the stable type causes instability, and its unstable type could determine the shadows of the black hole [21]. Also, in a new paper it was shown in a geometrical proof that ’any spherically symmetric space-time that is asymptotically flat and has a horizon must contain a photon sphere, a spherical photon shell’ [22]. Therefore, considering the above characteristics, it seems that the idea of the photon sphere can be used for a more precise classification of the space surrounding the ultra-compact structures.
@@
As, we have stated, the foundation of our work in this article is predicated on the necessity of the photon sphere for ultra-compact object structures. However, to approach the topological photon sphere historically, we commence our study of the photon sphere from the point where Cardoso and his colleagues have articulated: ”The mere observation of an unstable light ring is strong evidence for the existence of black holes”, or, ”The existence of a stable light ring is thus an unavoidable feature of any ultra-compact star” [23]. Then Cunha et al. showed during their studies that, in spherical symmetry, standard black holes, as well as other ultra-compact objects (that can be with or without event horizon) in general relativity, can have planar circular photon orbits. In such a way that the stable type causes instability, and its unstable type could determine the shadows of the black hole [22]. Finally, after Cunha et al. which extended their work to spherically symmetric 4D black holes[29],” Shao-Wen Wei” used this idea and” Duan” mapping, and extended the discussion of the photon ring to photon sphere with a topological approach. He stated that there exists at least one standard photon sphere outside the black hole not only in asymptotically flat space-time, but also in asymptotically AdS and dS space-time [24].
@@ 3 Explanation of topological charges based on properties of effective potential
In his paper, ’Wei’ demonstrated through the examination of winding numbers and the calculation of total charges for a model that whenever the TTC is -1, the structure will exhibit black hole behavior, and conversely, if the TTC is 0, the structure will be in the form of a naked singularity [24]. We have extended this work to show that in addition to the value of 0, the TTC value of +1 can also appear as a naked singularity [25]. In addition to utilizing the aforementioned analysis and topological charge analysis to classify the behavior of the studied structure and determine its parameter ranges, in this paper we also examine the geometric behavior of the effective potential (15), which shows precise alignment with the topological charges. A significant and distinct aspect of this potential, compared to traditional potentials used for studying photon spheres and null geodesics, is that it is independent of the energy and angular momentum of the incoming particle. This implies that this potential is an inherent property of the spacetime structure. Our observations in multiple cases [26, 27, 28] indicate that when the studied system exhibits a topological charge of -1, it always possesses the event horizon and displays black hole behavior. In such a scenario, the potential function reveals a dominant maximum outside the event horizon, indicating instability at the photon sphere location. Conversely, when the ultra-compact object under study exhibits a topological charge of 0 or +1, the system generally lacks an event horizon and takes the form of a naked singularity. In this case, the potential function reveals a dominant minimum alongside the maximum, indicating the presence of a stable photon sphere. It is worth noting an important and intriguing point here. From a classical perspective, the existence of a minimum energy state (stable condition) alongside a maximum energy state (unstable condition) implies that the system will naturally and spontaneously move towards the stable minimum. Consequently, it appears that the emergence of a stable photon sphere in proximity to an unstable photon sphere will cause any photon capable of leaving the unstable maximum to be absorbed by the nearest stable minimum. This suggests that the space-time in question should be considered ”off” from the perspective of observation and information reception.
Before dealing with the main models and for a better understanding of each of the topological states stated above, we will see an example separately below.
@@
As observed in Fig (1a), this black hole possesses a single zero point outside its event horizon, resulting in a TTC of -1, in accordance with reference [24]. In Fig (1b), we have plotted the potential H(r). It is evident that this potential has a global maximum, which upon closer inspection of Fig (1c), is precisely located at the zero point or in other words, at the same location as the photon sphere. Given that this point represents a global maximum, any photon trapped within this potential will abandon this peak at the slightest disturbance, in pursuit of a state with lower energy. This signifies that we possess an unstable photon sphere, precisely the phenomenon of observational interest.
Figure 1: Fig (1a):The normal vector field $n$ in the $(r-\theta)$ plane. The photon sphere is located at $(r,\theta)=(3.002309062,1.57)$ with respect to $(g=-0.21834,m=1,l=1)$ , (1b): the topological potential H(r) for regular Hayward AdS black hole model
@@
From an energy perspective, these two states occur when one or more local or global minima appear in the studied spacetime. For the topological charge 0 case, as observed in Figure (2b), a minimum—a stable photon sphere—at ( r = 1.389900878 ) can be clearly seen adjacent to a maximum—an unstable photon sphere. It is important to note a few points here. Firstly, this state typically occurs when, from the metric function’s viewpoint, we are in a rootless region, without a horizon, or in a space termed as a naked singularity. In other words, although these minima have always existed, even in the black hole state, but their influence on the final outcome only becomes apparent when the event horizon has disappeared, which for an ultra-compact object equates to a naked singularity. The second point to consider is whether the structural influence in the form of a naked singularity is unlimited. The answer depends on the model under study; sometimes, parametric conditions arise where neither the metric function nor the potential function has a solution for the studied interval. In other words, beyond that parametric range, neither the metric function nor the potential function practically provides a solution for the space, which we refer to as the forbidden region.
@@
In Fig(3), we encounter a scenario with a total topological charge (TTC) of +1.In this case, more than one minimum appears in the studied space-time.
@@ 4 Effective potential and topological photon sphere as a tool for classification
Before delving into the selected models, it is beneficial to briefly compare the advantages and disadvantages of this method relative to the traditional approach. In the traditional approach to studying photon spheres, it is essential to extract the Lagrangian from the action and subsequently construct the Hamiltonian. Once the Hamiltonian is determined, the effective potential can be constructed, which allows for the study of the photon sphere. This potential is a function of the energy and angular momentum of the particle.
@@
One notable advantage of this method is the direct use of the metric function to construct the potential, bypassing all the aforementioned steps. Additionally, the resulting potential is independent of the parameters of the incoming particle and is solely a function of the surrounding spacetime geometry. Using the equatorial plane for the vector field ($\phi$), similar to the Poincaré plane, and the resulting dimensional reduction, is another significant advantage. Ultimately, the ability to study a broader range of the space surrounding an ultra-compact object is a fundamental aspect that, for the first time, leads to the classification of different spacetime regions based on this method. Finally, this method allows for the introduction of a practical concept, which we term the possible radius limit for the photon sphere.
@@
$R_{PLPS}$: Considering the permissible parameter range to maintain the black hole structure, we define the minimum or maximum possible radius for the appearance of an unstable photon sphere as $R_{PLPS}$.
@@
However, there are fundamental drawbacks that must be considered. Since the vector field ($\phi$) must be separable with respect to spatial coordinates, only diagonal metrics with spherical symmetry can be used. Given that most models of interest have spherical structures and the rest can often be expressed diagonally using the C-metric, this limitation can be somewhat overlooked. Another issue is that since the Poincaré plane is used for dimensional reduction, this method may not be applicable to models with fewer than four dimensions. Finally, as this is a nascent approach, there may be weaknesses that become apparent over time. Interestingly, the simplifications, computational power, and novelty of this method and mathematical model for studying symmetric potential functions have garnered such attention that the mathematical structure of the model has been extended to thermodynamics. Consequently, extensive studies based on the mathematical model of this method have been conducted on the phase transitions of black hole models [33, 34, 35, 36, 37, 38, 39, 40, 41].
@@ 4.1 Photon Sphere and perfect fluid dark matter black hole(PFDM)
In the first case, considering the chosen value of the parameter $\lambda=0.62$ and the presence of the event horizon ($r_{h}=2.97163$) Fig (4), we observe the appearance of a photon sphere outside the event horizon. This photon sphere possesses a topological charge of -1 Fig (5a) and, as evident in Fig (5b), represents an energy maximum. Consequently, in this case, we are confronted with a black hole that contains an unstable photon sphere. However, in the second case, our choice of $\lambda=2.15$ results in a metric function without roots, which is clearly visible in Fig (4). The gravitational structure displays two photon sphere with charges of -1 and +1, indicating a space-time with a TTC of zero, Fig (6a). Energy-wise, this situation is equivalent to the emergence of both a minimum and a maximum, as shown in Fig (6b). Given these conditions, in this scenario, the structure manifests itself in the form of a naked singularity.
Now we have to answer the question that for what values of the parameter $\lambda$ the space will be in the form of a black hole or naked singularity? We have shown the answers to these questions in Table 1. Here, it is necessary to note that we have focused on the domain of $\lambda$ and assigned an arbitrary value to M. Obviously, changing these values can change our parameter range.